OpenStudy (anonymous):

I WILL GIVE MEDALS What is the slope of the graph of 4x - 8y = 11? Option A: -2 Option B: - ½ Option C: ½ Option D: 2

OpenStudy (anonymous):

I think it is C: 1/2

HanAkoSolo (jamierox4ev3r):

the key here is to get this equation in y=mx+b format So solve for y.

HanAkoSolo (jamierox4ev3r):

4x-8y=11 -4x -4x --------- -8y = -4x+11 \-8 \-8 ----------- y= \(\large\frac{1}{2}\)x - \(\large\frac{11}{8}\)

HanAkoSolo (jamierox4ev3r):

Do you understand how I solved for y? @LuJe

OpenStudy (anonymous):

Ya kinda

HanAkoSolo (jamierox4ev3r):

And you said that you think it's C :) That's correct, nice work

OpenStudy (anonymous):

Okay thanks! Could you help me with more @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

Sure